OpenStudy (anonymous):

Which type of triangle is formed by joining the vertices A(-3, 6), B(2, 1), and C(9, 5)?

OpenStudy (anonymous):

Get the length of each side by subtracting a-b = (x,y) square rout of x^2 + y^2. then if the longest side say a a^2< b^2 +c^2 then it is acute a^2 = b^2 + c^2 then it is right angle a^2 > b^2 + c^2 then it is obtuse angle